If the perimeter of a rectangle and square are equal and the ratio of the two adjacent sides of the rectangle is 1:2, then the ratio of area of the rectangle and that of the square is
Answer by LinnW(1048)

If width of the rectangle = 1
then its perimeter = 1 + 2 + 1 + 2 = 6
A square with perimeter 6 , 4s = 6
s = 1.5
The area of a square with side length 1.5 is
1.5*1.5 = 2.25
The area of a rectangle with width 1 and length 2
1*2 = 2
So the ratio of areas is 2/2.25 = 8/9 or 8:9